# Some characterizations of weakly uniformly perfect sets

### v5

@@ Abstract.
In this paper, the concept of weakly uniform perfectness is considered. As an analogue of the theory of uniform perfectness, we obtain the relationships between weakly uniform perfectness and Bergman kernel, Poincaré metric, local harmonic measure and Hausdorff content. In particular, for a bounded domain $\Omega\subset\mathbb{C}$, we show that the uniform perfectness of $\partial\Omega$ is equivalent to $K_{\Omega}(z)\gtrsim\delta(z)^{-2}$, where $K_{\Omega}(z)$ is the Bergman kernel of $\Omega$ and $\delta(z)$ denotes the boundary distance.
@@ 1. Introduction
The concept of uniform perfection is recognized to possess numerous equivalent characterizations (see,e.g., [2, 9, 19, 26, 30, 31]). Below, we enumerate a few. The uniform perfectness of a closed set $K$ is equivalent to the following statements:
@@
$K_{\Omega}(z)\gtrsim\delta(z)^{-2},$
@@
$b_{\Omega}(z)\gtrsim\delta(z)^{-1},$
where $\Omega$ is assumed to be a domain in $\mathbb{C}$ with $\partial\Omega=K$, and $\delta(z)$ denotes the distance of $z$ to $\partial\Omega$;
$(iii)$ the Poincaré metric $\rho(z)|\mathrm{d}z|$ of $\Omega$ satisfies $\rho(z)\gtrsim\delta(z)^{-1}$, where the domain $\Omega$ is as above;
$(iv)$ $\Omega$ is uniformly $\Delta$-regular, where the domain $\Omega$ is as above;
$(v)$ there exist constants $\alpha,A>0$, such that for every $a\in K$ and $r\in\left(0,\frac{\mathrm{diam}(\Omega)}{2}\right)$, we have
@@
In [34], a type of domains in $\mathbb{C}$ is considered, whose properties are analogous to domains with uniformly perfect boundary. In this paper, a closed set $K\subset\mathbb{C}$ is called $h$-uniformly perfect, if there exists a constant $r_{0}>0$ such that
@@
for every $a\in K$ and $r\in(0,r_{0})$, where $h$ is an increasing function on $[0,r_{0})$ with $h(r)\leq r$. In particular, if $K$ is bounded and $h$-uniformly perfect with $h(r)=cr\,(0<c<1)$, then $K$ is uniformly perfect.
We hope to delve deeper into this concept, but there are still some difficulties to consider the general situations, so we focus on some special cases in this paper. Let $h_{1,\alpha}(r)=r^{\alpha}$ and $h_{2,\beta}(r)=r(-\log r)^{-\beta}$ for some $\alpha>1$ and $\beta>0$. As in [34], we say $E\subset\mathbb{C}$ satisfies the condition $(U)_{1,\alpha}$ or $(U)_{2,\beta}$ if it is $Ch_{1,\alpha}$ or $Ch_{2,\beta}$-uniformly perfect for some $C>0$, respectively.
A generalization of statement (i) has already been obtained in [34]. It is also shown in [34] that if the boundary of $\Omega\subset\mathbb{C}$ satisfies $(U)_{1,\alpha}$ with $21<\alpha<$ 2 or $(U)_{2,\beta}$ with $\beta>0$, then we have
$K_{\Omega}(z)\gtrsim\frac{1}{\delta(z)^{2}\log\frac{1}{\delta(z)}},\,\,\,z\to\partial\Omega$
@@
$K_{\Omega}(z)\gtrsim\frac{1}{\delta(z)^{2}\log\log\frac{1}{\delta(z)}},\,\,\,z\to\partial\Omega,$
respectively. Here we shall prove the following theorem:
@@
$(1)$ If $\Omega$ is bounded and $K_{\Omega}(z)\gtrsim\delta(z)^{-2}$, then $\partial\Omega$ is uniformly perfect.
$(2)$If $K_{\Omega}(z)\geq\frac{M}{\delta(z)^{2}\log\frac{1}{\delta(z)}}$ for some $M>0$, then there exists some $\alpha>1$ such that $\partial\Omega$ satisfies $(U)_{1,\alpha}$. In particular, if $M>\frac{1}{2\pi}$, then $\alpha\in(1,2)$.
$(3)$ If $K_{\Omega}(z)\gtrsim\frac{1}{\delta(z)^{2}\log\log\frac{1}{\delta(z)}}$, then there exists some $\beta>0$ such that $\partial\Omega$ satisfies $(U)_{2,\beta}$.
In particular, the uniform perfectness of $\partial\Omega$ is equivalent to that $K_{\Omega}(z)\asymp\delta(z)^{-2}$ if $\Omega$ is bounded, which simplifies the statement $(ii)$. But it is still unknown whether the uniform perfectness of $\partial\Omega$ is equivalent to that $b_{\Omega}(z)\asymp\delta(z)^{-1}$.
@@
$(1)$ If $\rho_{\Omega}(z)\geq\frac{M}{\delta(z)\log\frac{1}{\delta(z)}}$ with $M>1$, then $\partial\Omega$ satisfies $(U)_{1,\alpha}$ for some $\alpha>1$.
$(2)$ $\rho_{\Omega}(z)\gtrsim\frac{1}{\delta(z)\log\log\frac{1}{\delta(z)}}\Leftrightarrow\partial\Omega$ satisfies $(U)_{2,\beta}$ for some $\beta>0$.
The condition $M>1$ in $(1)$ is sharp, since $\rho_{\mathbb{D}-\{0\}}(z)=\frac{1}{|z|\log\frac{1}{|z|}}$. The proof of Theorem 1.2 is analogous to that of Theorem 1.1, which is in fact inspired by Beardon-Pommerenke [4].
@@
for every $a\in\partial\Omega,r>0$ and $z\in\partial D(a,r/2)$, where $w_{x,r}$ denotes the harmonic measure of $\Omega\cap\partial D(a,r)$ in the domain $\Omega\cap D(a,r)$. Analogously, we say that a domain $\Omega\subset\mathbb{C}$ satisfies the condition $(\Delta)_{h}$, if there exist $\varepsilon\in(0,1)$ and $r_{0}>0$ such that
@@
for every $a\in\partial\Omega,r\in(0,r_{0})$ and $z\in\Omega\cap\partial D(a,h(r))$. For $\alpha>1$ and $\beta>0$, let $h_{1,\alpha}(t)=t^{\alpha},h_{2,\beta}(t)=t\left(\log\frac{1}{t}\right)^{-\beta}$. A domain $\Omega$ is said to satisfy $(\Delta)_{1,\alpha}$ or $(\Delta)_{2,\beta}$ if there exists a $C>0$ such that $\Omega$ satisfies $(\Delta)_{Ch_{1,\alpha}}$ or $(\Delta)_{Ch_{2,\beta}}$, respectively. We obtain
@@
$(1)$ If $\partial\Omega$ satisfies $(U)_{1,\alpha}$ for some $\alpha\in(1,2)$, then there exists a constant $\alpha^{\prime}>0$ such that $\Omega$ satisfies $(\Delta)_{1,\alpha^{\prime}}$. Conversely, if $\Omega$ satisfies $(\Delta)_{1,\alpha^{\prime}}$ for some $\alpha^{\prime}>0$, then there exists a constant $\alpha>1$ such that $\partial\Omega$ satisfies $(U)_{1,\alpha}$.
$(2)$ $\partial\Omega$ satisfies $(U)_{2,\beta}$ for some $\beta>0$Leftrightarrow $\Omega$ satisfies $(\Delta)_{2,\beta^{\prime}}$ for some $\beta^{\prime}>0$.
Now we focus on the statement (v), which implies that the Hausdorff dimension of a uniformly perfect set is always positive. However, for a closed set $E\subset\mathbb{C}$ only meets the condition $(U)_{1,\alpha}$ or $(U)_{2,\beta}$, it may not be detected by the classical Hausdorff dimension (see Appendix). Here we provide an analogue of statement (v):
@@
$(1)$ $E$ satisfies $(U)_{1,\alpha}$ for some $\alpha>1\Leftrightarrow$ there exist positive constants $A,C,r_{0}$ and $\gamma$ such that
@@
on $(0,2r_{0}).$
$(2)$ $E$ satisfies $(U)_{2,\beta}$ for some $\beta>0\Leftrightarrow$ there exist positive constants $A,r_{0}$ and $\eta$ such that
@@
is an orthnormal basis of $A^{2}(R)$, the Bergman kernel of $A^{2}(R)$ can be represented by
$K_{R}(z,w)=\frac{1}{2\pi z\overline{w}\log\frac{1}{s}}+\frac{1}{\pi z\overline{w}}\sum_{n\in\mathbb{Z}-\{0\}}\frac{n(z\overline{w})^{n}}{1-s^{2n}}.$
Let $z\in R$ with $|z|=s^{t}$, where $s,t\in(0,\frac{1}{2}]$, then
(2.1) $K_{R}(z)$ $=$ $\frac{1}{2\pi s^{2t}\log\frac{1}{s}}+\frac{1}{\pi s^{2t}}\left(\sum_{n=1}^{+\infty}\frac{ns^{2nt}}{1-s^{2n}}+\sum_{n=1}^{+\infty}\frac{(-n)s^{-2nt}}{1-s^{-2n}}\right)$
$\leq$ $\frac{1}{2\pi s^{2t}\log\frac{1}{s}}+\frac{1}{\pi s^{2t}}\sum_{n=1}^{\infty}\frac{n(s^{2nt}+s^{2n-2nt})}{1-s^{2n}}$
$\leq$ $\frac{1}{2\pi s^{2t}\log\frac{1}{s}}+\frac{2}{\pi s^{2t}}\sum_{n=1}^{+\infty}\frac{ns^{2nt}}{1-s^{2n}}$
$\leq$ $\frac{1}{2\pi s^{2t}\log\frac{1}{s}}+\frac{8}{3\pi s^{2t}}\sum_{n=1}^{+\infty}ns^{2nt}$
$=$ $\frac{1}{2\pi s^{2t}\log\frac{1}{s}}+\frac{8}{3\pi s^{2t}}\cdot\frac{s^{2t}}{(1-s^{2t})^{2}}$
$\leq$ $\frac{1}{2\pi s^{2t}\log\frac{1}{s}}+\frac{8}{3\pi}\cdot\frac{1}{(1-\frac{1}{2^{2t}})^{2}}$
$=$ $\frac{1}{2\pi s^{2t}\log\frac{1}{s}}+\mathrm{const}(t),$
where $\mathrm{const}(t)$ denotes a constant which only depends on $t$.In particular, if $t=\frac{1}{2}$, then $K_{R}(z)\leq\frac{1}{2\pi s\log\frac{1}{s}}+\frac{32}{3\pi}.$
@@
$(1)$ Notice that $\partial\Omega$ is uniformly perfect is equivalent to that there are constants $C\in(0,1)$ and $r_{0}>0$ such that
@@
We may take $C<\frac{1}{4},r_{0}<\mathrm{diam}(\Omega)/2$. Denote $A=\{z\in\mathbb{C};Cr<|z-a|<r\}$. Notice that $r_{0}<\mathrm{diam}(\Omega),$ if $\overline{A}\subset(\Omega^{c})^{o}$, then $\Omega$ cannot be connected, which implies $\overline{A}\subset\Omega$.
@@
Since $\delta(z_{0})\leq|z_{0}-a|=\sqrt{C}\cdot r$, we have
$K_{A}(z_{0})\leq\left(\frac{1}{2\pi C\log\frac{1}{C}}+\frac{32}{3\pi}\right)\cdot\frac{C}{\delta(z_{0})^{2}}.$
On the other hand, $A$ is a subdomain of $\Omega$, so
$K_{A}(z_{0})\geq K_{\Omega}(z_{0})\geq M\cdot\delta(z_{0})^{-2}$
for some $M>0$. Thus, for each $C\in(0,\frac{1}{4})$ and $r_{0}>0$, there exists a point $z_{0}\in\Omega$ with $\delta(z_{0})<r_{0}$, such that
@@
But this is impossible, since $C$ can be sufficiently small. Thus $\partial\Omega$ must be uniformly perfect.
$(2)$ Let $t\in(0,\frac{1}{2}]$. Suppose on the contrary that $\partial\Omega$ does not satisfy $(U)_{1,\alpha}$ for every $\alpha\in(1,2)$. Then for every $C>0$ and $r_{0}>0$, there exist $a\in\partial\Omega$ and $r\in(0,r_{0})$, such that
@@
Take $z_{0}\in A$ with $|z_{0}-a|=C^{t}r^{t(\alpha-1)+1}$. Write $z_{1}=T(z_{0})$, then $|z_{1}|=(Cr^{\alpha-1})^{t}$. Thus (2.1) implies
@@
$\delta(z_{0})\geq|z_{0}-a|-Cr^{\alpha}\geq\frac{C^{t}}{2}r^{t(\alpha-1)+1}$
for $r_{0}$ sufficiently small, we have
(2.3) $\left(\frac{\delta(z_{0})}{C^{t}}\right)^{\frac{1}{t(\alpha-1)+1}}\leq r\leq\left(\frac{2\delta(z_{0})}{C^{t}}\right)^{\frac{1}{t(\alpha-1)+1}}.$
@@
$K_{A}(z_{0})$$\leq$ $\mathrm{const}(t)\left(\frac{C^{t}}{\delta(z_{0})}\right)^{\frac{2}{t(\alpha-1)+1}}+\frac{1}{2\pi C^{2t}r^{2t(\alpha-1)+2}\left((\alpha-1)\log\frac{1}{r}-\log C\right)}$
$\leq$ $\mathrm{const}(t,C)\delta(z_{0})^{-\frac{2}{t(\alpha-1)+1}}+\frac{t(\alpha-1)+1}{2\pi(\alpha-1)\delta(z_{0})^{2}}\cdot\frac{1}{\left(\log\frac{1}{\delta(z_{0})}+\mathrm{const}(t,C)\right)}$
@@
$K_{A}(z_{0})\geq K_{\Omega}(z_{0})\geq M\cdot\frac{1}{\delta(z_{0})^{2}\log\frac{1}{\delta(z_{0})}},$
where $M>0$. Thus, for every $t\in(0,\frac{1}{2}],\alpha\in(1,\infty),C>0$ and $r_{0}>0$, there is a $z_{0}\in\Omega$ with $\delta(z_{0})\leq r_{0}$, such that
$0<M\leq const(t,C)\delta(z_{0})^{2-\frac{2}{t(\alpha-1)+1}}\log\frac{1}{\delta(z_{0})}+\frac{t(\alpha-1)+1}{2\pi(\alpha-1)}\cdot\frac{\log\frac{1}{\delta(z_{0})}}{\log\frac{1}{\delta(z_{0})}+const(t,C)}.$
Let $C$ be fixed and $r_{0}\to 0$, then
@@
$(3)$ Suppose on the contrary that $\partial\Omega$ does not satisfies $(U)_{2,\beta}$ for every $\beta>0$. Then for every $C>0$ and $r_{0}>0$, there are $a\in\partial\Omega$ and $r\in(0,r_{0})$, such that
@@
(2.5) $\delta(z_{0})\leq|z_{0}-a|=\sqrt{C}r\left(\log\frac{1}{r}\right)^{-\beta/2}$
@@
(2.6) $\delta(z_{0})\geq|z_{0}-a|-Cr\left(\log\frac{1}{r}\right)^{-\beta}\geq\frac{\sqrt{C}}{2}r\left(\log\frac{1}{r}\right)^{\beta/2}.$
From the proof of Lemma 5.2 in [34], it is easy to see that if
@@
(2.7) $r\leq\frac{2}{\sqrt{C}}\delta(z_{0})\left(\log\frac{1}{\delta(z_{0})}\right)^{\beta/2}.$
By (2.5), (2.6), (2.7) and (2.4), we have
$K_{A}(z_{0})\leq I\cdot\frac{1}{\delta(z_{0})^{2}},$

### v6

@@ Abstract.
In this paper, the concept of weakly uniform perfectness is considered. As an analogue of the theory of uniform perfectness, we establish the relationships between weakly uniform perfectness and Bergman kernel, Poincaré metric, local harmonic measure and Hausdorff content. In particular, for a bounded domain $\Omega\subset\mathbb{C}$, we show that the uniform perfectness of $\partial\Omega$ is equivalent to $K_{\Omega}(z)\gtrsim\delta_{\Omega}(z)^{-2}$, where $K_{\Omega}(z)$ is the Bergman kernel of $\Omega$ and $\delta_{\Omega}(z)$ denotes the boundary distance.
@@ 1. Introduction
The concept of uniform perfection is known to have numerous equivalent characterizations (see,e.g., [2, 9, 19, 26, 30, 31]). Below, we list a few: the uniform perfectness of a closed set $K$ is equivalent to the following statements:
@@
$K_{\Omega}(z)\gtrsim\delta_{\Omega}(z)^{-2},$
@@
$b_{\Omega}(z)\gtrsim\delta_{\Omega}(z)^{-1},$
where $\Omega$ is assumed to be a domain in $\mathbb{C}$ with $\partial\Omega=K$, and $\delta_{\Omega}(z)$ denotes the distance of $z$ to $\partial\Omega$;
$(iii)$ the Poincaré metric $\rho_{\Omega}(z)|\mathrm{d}z|$ of $\Omega$ satisfies $\rho_{\Omega}(z)\gtrsim\delta_{\Omega}(z)^{-1}$, where the domain $\Omega$ is assumed as above;
$(iv)$ $\Omega$ is uniformly $\Delta$-regular, where the domain $\Omega$ is assumed as above;
$(v)$ there exist constants $\alpha,A>0$ such that for every $a\in K$ and $r\in\left(0,\frac{\mathrm{diam}(\Omega)}{2}\right)$, we have
@@
In [34], a type of domains in $\mathbb{C}$ is studied, whose properties are analogous to those of domains with uniformly perfect boundary. In this paper, a closed set $K\subset\mathbb{C}$ is called $h$-uniformly perfect if there exists a constant $r_{0}>0$ such that
@@
for every $a\in K$ and $r\in(0,r_{0})$, where $h$ is an increasing function on $0,r_{0})$ satisfying $h(r)\leq r$. In particular, if $K$ is bounded and $h$-uniformly perfect with h(r)=cr(0<c<1)h(r)=cr\,(0<c<1), then $K$ is uniformly perfect.
We hope to explore this concept more deeply; however, due to certain difficulties in addressing the general situations, we focus on some special cases in this paper. Let $h_{1,\alpha}(r)=r^{\alpha}$ and $h_{2,\beta}(r)=r(-\log r)^{-\beta}$ for some $\alpha>1$ and $\beta>0$. Following [34], we say that a set $E\subset\mathbb{C}$ satisfies condition $(U)_{1,\alpha}$ or $(U)_{2,\beta}$ if it is $Ch_{1,\alpha}$-uniformly perfect or $Ch_{2,\beta}$-uniformly perfect for some $C>0$, respectively.
A generalization of statement (i) has already been obtained in [34]. It is also shown in [34] that if the boundary of $\Omega\subset\mathbb{C}$ satisfies $(U)_{1,\alpha}$ with $1<\alpha<2$ or $(U)_{2,\beta}$ with $\beta>0$, then we have
$K_{\Omega}(z)\gtrsim\frac{1}{\delta_{\Omega}(z)^{2}\log\frac{1}{\delta_{\Omega}(z)}},\,\,\,z\to\partial\Omega$
@@
$K_{\Omega}(z)\gtrsim\frac{1}{\delta_{\Omega}(z)^{2}\log\log\frac{1}{\delta_{\Omega}(z)}},\,\,\,z\to\partial\Omega,$
respectively. Here, we shall prove the following theorem:
@@
$(1)$ If $\Omega$ is bounded and $K_{\Omega}(z)\gtrsim\delta_{\Omega}(z)^{-2}$, then $\partial\Omega$ is uniformly perfect.
(2)(2)If $K_{\Omega}(z)\geq\frac{M}{\delta_{\Omega}(z)^{2}\log\frac{1}{\delta_{\Omega}(z)}}$ for some $M>0$, then there exists some $\alpha>1$ such that $\partial\Omega$ satisfies condition $(U)_{1,\alpha}$. In particular, if $M>\frac{1}{2\pi}$, then $\alpha\in(1,2)$.
$(3)$ If $K_{\Omega}(z)\gtrsim\frac{1}{\delta_{\Omega}(z)^{2}\log\log\frac{1}{\delta_{\Omega}(z)}}$, then there exists $\beta>0$ such that $\partial\Omega$ satisfies condition $(U)_{2,\beta}$.
In particular, if $\Omega$ is bounded, the uniform perfectness of $\partial\Omega$ is equivalent to that $K_{\Omega}(z)\asymp\delta_{\Omega}(z)^{-2}$ , which simplifies the statement $(ii)$. However, it remains unknown whether the uniform perfectness of $\partial\Omega$ is also equivalent to $b_{\Omega}(z)\asymp\delta_{\Omega}(z)^{-1}$.
@@
$(1)$ If $\rho_{\Omega}(z)\geq\frac{M}{\delta_{\Omega}(z)\log\frac{1}{\delta_{\Omega}(z)}}$ for some $M>1$, then $\partial\Omega$ satisfies $(U)_{1,\alpha}$ for some $\alpha>1$.
$(2)$ $\rho_{\Omega}(z)\gtrsim\frac{1}{\delta_{\Omega}(z)\log\log\frac{1}{\delta_{\Omega}(z)}}\Leftrightarrow\partial\Omega$ satisfies $(U)_{2,\beta}$ for some $\beta>0$.
The condition $M>1$ in $(1)$ is sharp, since $\rho_{\mathbb{D}-\{0\}}(z)=\frac{1}{|z|\log\frac{1}{|z|}}$. The proof of Theorem 1.2 is analogous to that of Theorem 1.1, which is inspired by Beardon-Pommerenke [4].
@@
for every $a\in\partial\Omega,r\in(0,\mathrm{diam}(\Omega))$, and $z\in\partial D(a,r/2)$, where $w_{a,r}$ denotes the harmonic measure of $\Omega\cap\partial D(a,r)$ in the domain $\Omega\cap D(a,r)$. Analogously, we say that a domain $\Omega\subset\mathbb{C}$ satisfies condition $(\Delta)_{h}$, if there exist $\varepsilon\in(0,1)$ and $r_{0}>0$ such that
@@
for every $a\in\partial\Omega,r\in(0,r_{0})$, and $z\in\Omega\cap\partial D(a,h(r))$. For $\alpha>1$ and $\beta>0$, let $h_{1,\alpha}(t)=t^{\alpha}$, and $h_{2,\beta}(t)=t\left(\log\frac{1}{t}\right)^{-\beta}$. A domain $\Omega$ is said to satisfy $(\Delta)_{1,\alpha}$ or $(\Delta)_{2,\beta}$ if there exists a $C>0$ such that $\Omega$ satisfies condition $(\Delta)_{Ch_{1,\alpha}}$ or $(\Delta)_{Ch_{2,\beta}}$, respectively. We obtain
@@
$(1)$ If $\partial\Omega$ satisfies condition $(U)_{1,\alpha}$ for some $\alpha\in(1,2)$, then there exists a constant $\alpha^{\prime}>0$ such that $\Omega$ satisfies $(\Delta)_{1,\alpha^{\prime}}$. Conversely, if $\Omega$ satisfies $(\Delta)_{1,\alpha^{\prime}}$ for some $\alpha^{\prime}>0$, then there exists a constant $\alpha>1$ such that $\partial\Omega$ satisfies $(U)_{1,\alpha}$.
$(2)$ $\partial\Omega$ satisfies $(U)_{2,\beta}$ for some $\beta>0$ if and only if $\Omega$ satisfies $(\Delta)_{2,\beta^{\prime}}$ for some $\beta^{\prime}>0$.
Now we focus on the statement (v), which implies that the Hausdorff dimension of a uniformly perfect set is always positive. However, for a closed set $E\subset\mathbb{C}$ only meets condition $(U)_{1,\alpha}$ or $(U)_{2,\beta}$, it may not be detected by the classical Hausdorff dimension (see Appendix). Here we provide an analogue of statement (v):
@@
$(1)$ $E$ satisfies contion $(U)_{1,\alpha}$ for some $\alpha>1\Leftrightarrow$ there exist positive constants $A,C,r_{0}$ and $\gamma$ such that
@@
on $(0,2r_{0}),$ and $\Lambda_{g_{1,\gamma}}$ denotes the $g_{1,\gamma}$-Hausdorff content.
$(2)$ $E$ satisfies condition $(U)_{2,\beta}$ for some $\beta>0\Leftrightarrow$ there exist positive constants $A,r_{0}$ and $\eta$ such that
@@
forms an orthnormal basis of $A^{2}(R)$, the Bergman kernel of $A^{2}(R)$ can be represented by
$K_{R}(z,w)=\frac{1}{2\pi z\overline{w}\log\frac{1}{r}}+\frac{1}{\pi z\overline{w}}\sum_{n\in\mathbb{Z}-\{0\}}\frac{n(z\overline{w})^{n}}{1-r^{2n}}.$
Let $z\in R$ with $|z|=r^{t}$, where $r,t\in(0,\frac{1}{2}]$, then
(2.1) $K_{R}(z)$ $=$ $\frac{1}{2\pi r^{2t}\log\frac{1}{r}}+\frac{1}{\pi r^{2t}}\left(\sum_{n=1}^{+\infty}\frac{nr^{2nt}}{1-r^{2n}}+\sum_{n=1}^{+\infty}\frac{(-n)r^{-2nt}}{1-r^{-2n}}\right)$
$\leq$ $\frac{1}{2\pi r^{2t}\log\frac{1}{r}}+\frac{1}{\pi r^{2t}}\sum_{n=1}^{\infty}\frac{n(r^{2nt}+r^{2n-2nt})}{1-r^{2n}}$
$\leq$ $\frac{1}{2\pi r^{2t}\log\frac{1}{r}}+\frac{2}{\pi r^{2t}}\sum_{n=1}^{+\infty}\frac{nr^{2nt}}{1-r^{2n}}$
$\leq$ $\frac{1}{2\pi r^{2t}\log\frac{1}{r}}+\frac{8}{3\pi r^{2t}}\sum_{n=1}^{+\infty}nr^{2nt}$
$=$ $\frac{1}{2\pi r^{2t}\log\frac{1}{r}}+\frac{8}{3\pi r^{2t}}\cdot\frac{r^{2t}}{(1-r^{2t})^{2}}$
$\leq$ $\frac{1}{2\pi r^{2t}\log\frac{1}{r}}+\frac{8}{3\pi}\cdot\frac{1}{(1-\frac{1}{2^{2t}})^{2}}$
$=$ $\frac{1}{2\pi r^{2t}\log\frac{1}{r}}+\mathrm{const}(t),$
where $\mathrm{const}(t)$ denotes a constant that depends only on $t$. In particular, if $t=\frac{1}{2}$, then
$K_{R}(z)\leq\frac{1}{2\pi r\log\frac{1}{r}}+\frac{32}{3\pi}.$
@@
$(1)$ Notice that $\partial\Omega$ is uniformly perfect is equivalent to the existence of constants $C\in(0,1)$ and $r_{0}>0$ such that
@@
We may take $C<\frac{1}{4},r_{0}<\frac{\mathrm{diam}(\Omega)}{2}$. Denote $A=\{z\in\mathbb{C};Cr<|z-a|<r\}$. Notice that $r_{0}<\mathrm{diam}(\Omega),$ if $\overline{A}\subset(\Omega^{c})^{o}$, then $\Omega$ is not connected. Therefore, $\overline{A}\subset\Omega$.
@@
Since $\delta_{\Omega}(z_{0})\leq|z_{0}-a|=\sqrt{C}\cdot r$, we have
$K_{A}(z_{0})\leq\left(\frac{1}{2\pi C\log\frac{1}{C}}+\frac{32}{3\pi}\right)\cdot\frac{C}{\delta_{\Omega}(z_{0})^{2}}.$
On the other hand, since $A$ is a subdomain of $\Omega$, we have
$K_{A}(z_{0})\geq K_{\Omega}(z_{0})\geq M\cdot\delta_{\Omega}(z_{0})^{-2}$
for some constant $M>0$. Thus, for each $C\in(0,\frac{1}{4})$ and $r_{0}>0$, there exists a point $z_{0}\in\Omega$ with $\delta_{\Omega}(z_{0})<r_{0}$ such that
@@
But this leads to a contradiction, since $C$ can be taken sufficiently small. Therefore, $\partial\Omega$ must be uniformly perfect.
$(2)$ Let $t\in(0,\frac{1}{2}]$. Suppose, the contrary, that $\partial\Omega$ does not satisfy condition $(U)_{1,\alpha}$ for every $\alpha\in(1,2)$. Then for every $C>0$ and $r_{0}>0$, there exist $a\in\partial\Omega$ and $r\in(0,r_{0})$ such that
@@
Take $z_{0}\in A$ with $|z_{0}-a|=C^{t}r^{t(\alpha-1)+1}$. Write $z_{1}=T(z_{0})$, then $|z_{1}|=(Cr^{\alpha-1})^{t}$. Thus, (2.1) implies
@@
$\delta_{\Omega}(z_{0})\geq|z_{0}-a|-Cr^{\alpha}\geq\frac{C^{t}}{2}r^{t(\alpha-1)+1}$
for sufficiently small $r_{0}$, we have
(2.3) $\left(\frac{\delta_{\Omega}(z_{0})}{C^{t}}\right)^{\frac{1}{t(\alpha-1)+1}}\leq r\leq\left(\frac{2\delta_{\Omega}(z_{0})}{C^{t}}\right)^{\frac{1}{t(\alpha-1)+1}}.$
@@
$K_{A}(z_{0})$ $\leq$ $\mathrm{const}(t)\left(\frac{C^{t}}{\delta_{\Omega}(z_{0})}\right)^{\frac{2}{t(\alpha-1)+1}}+\frac{1}{2\pi C^{2t}r^{2t(\alpha-1)+2}\left((\alpha-1)\log\frac{1}{r}-\log C\right)}$
$\leq$ $\mathrm{const}(t,C)\delta_{\Omega}(z_{0})^{-\frac{2}{t(\alpha-1)+1}}+\frac{t(\alpha-1)+1}{2\pi(\alpha-1)\delta_{\Omega}(z_{0})^{2}}\cdot\frac{1}{\left(\log\frac{1}{\delta_{\Omega}(z_{0})}+\mathrm{const}(t,C,\alpha)\right)}$
@@
$K_{A}(z_{0})\geq K_{\Omega}(z_{0})\geq M\cdot\frac{1}{\delta_{\Omega}(z_{0})^{2}\log\frac{1}{\delta_{\Omega}(z_{0})}},$
where $M>0$. Thus, for every $t\in(0,\frac{1}{2}],\alpha\in(1,\infty),C>0$ and $r_{0}>0$, there exists a point $z_{0}\in\Omega$ with $\delta_{\Omega}(z_{0})\leq r_{0}$, such that
$0<M\leq\mathrm{const}(t,C)\delta_{\Omega}(z_{0})^{2-\frac{2}{t(\alpha-1)+1}}\log\frac{1}{\delta_{\Omega}(z_{0})}+\frac{t(\alpha-1)+1}{2\pi(\alpha-1)}\cdot\frac{\log\frac{1}{\delta_{\Omega}(z_{0})}}{\log\frac{1}{\delta_{\Omega}(z_{0})}+\mathrm{const}(t,C,\alpha)}.$
Let $C$ be fixed and let $r_{0}\to 0$, then
@@
$(3)$ Suppose on the contrary that $\partial\Omega$ does not satisfies condition $(U)_{2,\beta}$ for every $\beta>0$. Then, for every $C>0$ and $r_{0}>0$, there exist $a\in\partial\Omega$ and $r\in(0,r_{0})$ such that
@@
(2.5) $\delta_{\Omega}(z_{0})\leq|z_{0}-a|=\sqrt{C}r\left(\log\frac{1}{r}\right)^{-\beta/2}$
@@
(2.6) $\delta_{\Omega}(z_{0})\geq|z_{0}-a|-Cr\left(\log\frac{1}{r}\right)^{-\beta}\geq\frac{\sqrt{C}}{2}r\left(\log\frac{1}{r}\right)^{-\beta/2}.$
From the proof of Lemma 5.2 in [34], it is clear that if
@@
(2.7) $r\leq\frac{2}{\sqrt{C}}\delta_{\Omega}(z_{0})\left(\log\frac{1}{\delta_{\Omega}(z_{0})}\right)^{\beta/2}.$
By (2.5), (2.6), (2.7), and (2.4), we have
$K_{A}(z_{0})\leq I\cdot\frac{1}{\delta_{\Omega}(z_{0})^{2}},$